# A Survey on Knowledge Distillation of Large Language Models

### v3

@@ 5.5 Miscellaneous
The expansion of LLMs into various verticals beyond the ones previously discussed showcases their versatility and transformative potential across numerous industries and societal sectors. LLMs are being tailored to meet the specific needs and challenges of different domains, from legal and governmental to entertainment and beyond, providing sophisticated natural language understanding, generation, and decision-making capabilities.
Education. EduChat (Dan et al., 2023) is a large-scale language model-based chatbot system designed for the education domain. It aims to revolutionize intelligent education by providing personalized, fair, and compassionate support to teachers, students, and parents. Knowledge distillation is emphasized through the pre-training on an educational corpus and fine-tuning on custom instructions to activate education-specific functions like open question answering, essay assessment, and emotional support. EduChat demonstrates the importance of domain-specified knowledge distillation in enhancing the performance of LLMs within specific verticals, offering a significant contribution to intelligent education technology.
IT Operations. Owl (Guo et al., 2023b) is a specialized LLM tailored for IT operations, focusing on enhancing efficiency and analysis within this domain. The model leverages a unique Owl-Instruct dataset covering a broad range of IT-related information, employing a mixture-of-adapter strategy for efficient domain-specific tuning. This approach significantly improves IT operation tasks’ performance, highlighting the critical role of knowledge distillation in adapting general LLMs to specialized fields such as IT operations, thereby pushing forward the frontier in specialized AI applications within this sector.
@@ 6 Open Problems
Trustworthy Knowledge Distillation Trustworthiness in LLMs is paramount, encompassing attributes such as truthfulness, safety, fairness, robustness, privacy, and adherence to machine ethics Sun et al. (2024a). The rapid advancement of LLMs brings to the forefront concerns regarding their trustworthiness, stemming from their complex outputs, the biases present in vast training datasets, and the potential inclusion of private information. Current efforts in KD of LLMs primarily focus on distilling various skills from LLMs, with relatively little attention paid to trustworthiness aspects. Existing studies tend to concentrate on a subset of trustworthiness aspects, such as helpfulness, honesty, and harmlessness Bai et al. (2022a); Yang et al. (2024a); Cui et al. (2023a). Consequently, in the distillation process, student models may inherit issues related to trustworthiness from their teacher LLMs. As assessed in Sun et al. (2024a), smaller open-source LLMs generally fall short of their proprietary counterparts in trustworthiness metrics. Therefore, considering trustworthiness alongside the distillation of capabilities into student models is crucial. It is imperative that future research on KD not only enhances the capabilities of student models but also ensures that broader aspects of trustworthiness are meticulously addressed.
@@
Weak-to-strong Distillation. The concept of “weak-to-strong generalization” in LLMs (Burns et al., 2023) emphasizes the potential to leverage weak supervision to elicit the advanced capabilities of more powerful models. This approach challenges the traditional distillation paradigm by suggesting that even with limited or imperfect supervision, it is possible to enhance the performance of LLMs significantly. This necessitates exploring innovative strategies that enable weaker models to guide the learning process of stronger ones effectively, highlighting the importance of developing methods that can bridge the gap between these models. Such research could unlock new avenues for improving LLMs’ efficiency and effectiveness, making the pursuit of “weak-to-strong distillation” a crucial area for future investigations in this LLM era. Initially, Burns et al. (2023) investigate whether weak model supervision can unlock the full capabilities of much stronger models. Through experiments with pre-trained language models in the GPT-4 family across NLP, chess, and reward modeling tasks, it finds that finetuning strong models on weak labels leads to better performance than their weak supervisors, demonstrating weak-to-strong generalization. Then, Li et al. (2024g) introduce Superfiltering, a method that employs smaller, weaker models like GPT-2 to select high-quality data for fine-tuning larger, more capable models such as LLaMA2. This approach is rooted in discovering a strong consistency in evaluating instruction tuning data difficulty across models of varying sizes. More recently, Ji et al. (2024) introduce Aligner, a novel approach for aligning LLMs with human values and intentions by utilizing weak supervisory signals from smaller models to improve the performance of larger models. However, Burns et al. (2023) find that achieving the full capabilities of strong models requires more than naive finetuning, suggesting the need for further research in this area. Therefore, open questions still remain about 1) What are the theoretical and practical limits of weak-to-strong distillation? Can weak supervision reliably extract and enhance the full spectrum of capabilities in stronger models across all domains, or are there inherent limitations based on model architecture or task specificity? 2) How do we identify or design the optimal weak supervisors for distilling knowledge into stronger models? Is there a framework or criteria to predict which weak models would be most effective in guiding the learning process of more complex models for specific tasks? 3) To what extent are weak-to-strong distillation techniques transferable and scalable across different sizes and types of models? How can these methods be adapted to ensure efficacy and efficiency in distilling knowledge from very large models to significantly smaller ones, especially in resource-constrained environments?
@@ 7 Conclusion and Discussion
This survey has traversed the expansive domain of knowledge distillation applied to LLMs, shedding light on the myriad techniques, applications, and emerging challenges in this vibrant field. We have underscored the pivotal role of KD in democratizing access to the advanced capabilities of proprietary LLMs, thereby fostering a more equitable AI landscape. Through meticulous examination, we have highlighted how KD serves as a bridge, enabling resource-constrained entities to benefit from the profound advancements in LLMs without the prohibitive costs associated with training and deploying state-of-the-art models.
Our exploration delineates the multifaceted approaches to KD, ranging from algorithmic innovations and skill enhancement to domain-specific distillations. Each segment reveals the nuanced complexities and potentialities inherent in tailoring distilled models to emulate the sophisticated understandings and functionalities of their more cumbersome counterparts. Notably, the integration of data augmentation strategies within KD processes emerges as a critical lever for enhancing distillation in this LLM era, underscoring the synergistic potential between generating context-rich training data and the distillation endeavor.
As we project into the future, several avenues for research beckon. The evolving landscape of AI, marked by rapid advancements in model architectures and training methodologies, presents both challenges and opportunities for KD. The quest for more efficient, transparent, and ethical AI models necessitates continued innovation in KD techniques, especially those that can navigate the delicate balance between model fidelity, computational efficiency, and ethical considerations. Furthermore, the exploration of KD in nascent areas such as weak-to-strong generalization, self-alignment, multi-modal LLMs, real-time adaptation, and personalized AI services promises to expand the horizons of what distilled models can achieve.
Therefore, knowledge distillation of LLMs stands at a critical juncture, embodying the potential to significantly influence the trajectory of AI development and application. As this survey elucidates, the concerted efforts of the research community in pushing the boundaries of KD will be instrumental in realizing the vision of accessible, efficient, and responsible AI for all.
Legal Considerations for Using LLM Outputs: Importantly, it’s crucial to note the legal implications of utilizing LLM outputs, such as those from ChatGPT [^4] https://openai.com/policies/business-terms, Llama [^5] https://llama.meta.com/llama-downloads/, etc. We strongly advocate compliance with the terms of use specified by the model providers, such as the restrictions on developing competitive products, and so on.

### v4

@@ 5.5 Miscellaneous
Knowledge distillation of LLMs has vast potential across various verticals beyond the ones previously discussed, highlighting their versatility and transformative impact across different industries. For instance, in the education sector, EduChat Dan et al. (2023) exemplifies a chatbot system that provides tailored support to teachers, students, and parents. KD is central to its design, leveraging pre-training on educational data followed by fine-tuning with custom instructions to deliver capabilities such as essay evaluation and emotional support. Similarly, Owl Guo et al. (2023b), an LLM designed for IT operations, boosts operational efficiency using the Owl-Instruct dataset, which is distilled from ChatGPT. By applying a mixture-of-adapter strategy for domain-specific tuning, it enhances analysis and performance in IT-related tasks.
@@ 6 Open Problems
Trustworthy Knowledge Distillation Trustworthiness in LLMs is paramount, encompassing attributes such as truthfulness, safety, fairness, robustness, privacy, and adherence to machine ethics Sun et al. (2024a). The rapid advancement of LLMs brings to the forefront concerns regarding their trustworthiness, stemming from their complex outputs, the biases present in vast training datasets, and the potential inclusion of private information. Current efforts in KD of LLMs primarily focus on distilling various skills from LLMs, with relatively little attention paid to trustworthiness aspects. Existing studies tend to concentrate on a subset of trustworthiness aspects, such as helpfulness, honesty, and harmlessness Bai et al. (2022a); Yang et al. (2024); Cui et al. (2023a). Consequently, in the distillation process, student models may inherit issues related to trustworthiness from their teacher LLMs. As assessed in Sun et al. (2024a), smaller open-source LLMs generally fall short of their proprietary counterparts in trustworthiness metrics. Therefore, considering trustworthiness alongside the distillation of capabilities into student models is crucial. It is imperative that future research on KD not only enhances the capabilities of student models but also ensures that broader aspects of trustworthiness are meticulously addressed.
@@
Weak-to-strong Distillation. The concept of “weak-to-strong generalization” in LLMs (Burns et al., 2023) emphasizes the potential to leverage weak supervision to elicit the advanced capabilities of more powerful models. This approach challenges the traditional distillation paradigm by suggesting that even with limited or imperfect supervision, it is possible to enhance the performance of LLMs significantly. This necessitates exploring innovative strategies that enable weaker models to guide the learning process of stronger ones effectively, highlighting the importance of developing methods that can bridge the gap between these models. Such research could unlock new avenues for improving LLMs’ efficiency and effectiveness, making the pursuit of “weak-to-strong distillation” a crucial area for future investigations in this LLM era. Initially, Burns et al. (2023) investigates whether weak model supervision can unlock the full capabilities of much stronger models. Through experiments with pre-trained language models in the GPT-4 family across NLP, chess, and reward modeling tasks, it finds that finetuning strong models on weak labels leads to better performance than their weak supervisors, demonstrating weak-to-strong generalization. Then, Li et al. (2024g) introduce Superfiltering, a method that employs smaller, weaker models like GPT-2 to select high-quality data for fine-tuning larger, more capable models such as LLaMA2. This approach is rooted in discovering a strong consistency in evaluating instruction tuning data difficulty across models of varying sizes. More recently, Ji et al. (2024) introduce Aligner, a novel approach for aligning LLMs with human values and intentions by utilizing weak supervisory signals from smaller models to improve the performance of larger models. However, Burns et al. (2023) find that achieving the full capabilities of strong models requires more than naive finetuning, suggesting the need for further research in this area. Therefore, open questions still remain about 1) What are the theoretical and practical limits of weak-to-strong distillation? Can weak supervision reliably extract and enhance the full spectrum of capabilities in stronger models across all domains, or are there inherent limitations based on model architecture or task specificity? 2) How do we identify or design the optimal weak supervisors for distilling knowledge into stronger models? Is there a framework or criteria to predict which weak models would be most effective in guiding the learning process of more complex models for specific tasks? 3) To what extent are weak-to-strong distillation techniques transferable and scalable across different sizes and types of models? How can these methods be adapted to ensure efficacy and efficiency in distilling knowledge from very large models to significantly smaller ones, especially in resource-constrained environments?
@@ 7 Conclusion and Discussion
This survey has explored the diverse landscape of knowledge distillation for LLMs, highlighting key techniques, applications, and challenges. KD plays a crucial role in democratizing access to advanced LLM capabilities, providing cutting-edge advancements without the high costs of training and deployment. Our review emphasizes various KD approaches, from algorithmic innovations to skill enhancement and vertical distillation. Notably, data augmentation and synthesis within KD emerge as vital tools for improving distillation, revealing the powerful synergy between enriched training data and effective model distillation. As the AI landscape evolves, rapid advancements in model architectures and training methods present both challenges and research opportunities for KD of LLMs. Future innovation will need to focus on achieving efficiency, transparency, and ethics while maintaining model trustworthiness. Furthermore, promising areas such as weak-to-strong generalization, self-alignment, and multi-modal LLMs offer the potential to enhance the capabilities of distilled models. In conclusion, the KD of LLMs is set to play a pivotal role in the future of AI research. As highlighted in this survey, sustained research efforts will be critical in developing accessible, efficient, and responsible AI for all. Importantly, when conducting KD of LLMs like ChatGPT or Llama, it’s essential to comply with the model providers’ terms [^4] OpenAI Business Terms: https://openai.com/policies/business-terms, such as the restrictions on developing competitive products.